Monthly Archives: May 2014

The week of May 26

Slightly shorter week since Monday is a holiday. The main goal for the week is the 97 point test. This is a mapping of the tracker surface such that we hope to generate a series of trajectories that map out a sphere above the primary. This sphere should be centered on the geometric center of the upper hex and have the same shape as the primary mirror. This will require that our Telescope Control System (TCS) is up and running. This will be the first test of the TCS systems and we should expect the normal number of communication protocol problems and software glitches. This 97 point test will be a significant milestone for the WFU project.

The duty Telescope Operator and the mirror team will continue to get the primary mirror in shape by aligning the mirrors every night and adjusting actuators and sensors every morning.

The week of May 19

In the past week we began several new projects at the HET. One is the Remote Thermal Area. This week that project continues with the contractors from Veliz Construction and BP locating all of our buried water, glycol and lightening protection lines around the back of the facility and starting the trench work for the new lines. This work should continue all week.

The other project is the sealing of the dome louvers. Over the years the brass seals on these louvers have blown off and we have had significant ingress of dust and rain under high winds. We have identified a new adhesive that should keeps those seals on over a much longer time scale. One of our TOs is working on the installation of these seals and based on last weeks work she finds that it takes a full day to install a seal all the way around each louver and there are 4 louvers in each bank. She hopes to be finished by the end of July.

Mirror alignment was begun last week. This is a fairly slow process and we have captured about 70 mirrors into our array and the rest require some by hand alignment, which will be done this week. This is moving along steadily and should be ready by the deadline to stay to schedule… the 2nd week of June.

Last week we made considerable progress in getting the track to run consistently. The source of the metal shavings was located: with the additional weight of the tracker the encoder track deflected and the encoder trolley had to be realigned. We will continue to fine tune the encoder head position as we add more weight to the tracker.

The Constant Force Drive is now operational and can move the track at will in the Y direction. We have repaired the burned out cards but we have not identified the source of that damage. This week we have a few odds and ends to prepare for next week’s 97 point tracking test.

The week of May 12

In the past week we made some progress with the Constant Force Drive (CFD). We found that one of the units of the Current to Voltage board had burned out. We also found that all of our spares have had a similar problem. We ordered an industrial drop-in part which does not quite fit the form factor for the electronics rack but did allow us to continue commissioning. We still have not determined why the boards were burned out but we will try to refurbish the parts and replace them as time allows. We also have one of our Telescope Operators now running the tracker. On a less positive note we found some metal fillings in the lower X encoder. This may simply be the race finding a new home under the weight of the tracker but we need some additional testing to determine if further damage may be occurring.

The mirror team was able to get all of the edge sensors back on the mirrors but we have run into significant communication problems with the CCAS alignment tower that did not allow us to fine align any of the mirrors in the array.

This week will be taken up with getting the communication problems with the CCAS alignment tower sorted out, investigating the problems with the lower X encoder, and trouble shooting the Current to Voltage boards.

We will also have contractors from Veliz Construction on site to start our big Remote Thermal Area project which actually includes a number of small projects including: Ducting heat from the Krabbenhut away from the telescope, ducting away and diluting acid vapor from the strip and wash room, installing the remote air conditioning and glycol chilling units, and installing the huge liquid nitrogen tank for the VIRUS instrument.

The week of May 5

It has been a few weeks since the last blog post. The blogger was out of town…

In the past few weeks there has been considerable progress. The hexapod assembly and PFIP have been mounted on the tracker. Many of the tests have pointed to significant communications troubles with different circuit boards, broken wires etc.. Many of these have been resolved and we have fairly stable X motion of the tracker but the Constant Force Drive (CFD) has been not been behaving well so that has limited the use of the tracker in Y. Since most of tracker cabling work and repainting of the upper hex are complete we have been able to remove the safety net above the primary mirror and begin installing the mirrors. In the past week the mirror team removed the net and installed 40 recoated mirrors.

In the rest of this week the mirror team will be installing the last mirrors (leaving a few holes in the array so that we can install test spots for the laser tracker). Once the mirrors sensors are in place we will begin aligning (stacking) all of the mirrors with night operations. This will be the first night time operations for the Telescope Operators since August 2013. Once the mirrors are aligned the mirror team will piston the array. This work will likely take the rest of the week.

In parallel, we hope to have some new circuits in place later this week which will allow the tracker team to test the CFD and Y motions again. Once those are working then they can test the hexapod motion.

Other work that will be done this week in parallel are the work on the purge air lines, installing mounting plates for future testing equipment and a punch-list of software and hardware work.